OpenStudy (anonymous):

Solve the ODE x^2y''-xy'+=ln(x) I will rewrite the equation in the editor and show what I have done so far below.

OpenStudy (anonymous):

\[x^{2}y''-xy'+y=\ln(x)\] So I solved the homogeneous part as follows: \[r(r-1)-r+1=0 \rightarrow r^2-2r+1=0 \rightarrow (r-1)^2=0 \rightarrow r _1=r_2=1\] Thus \[y_H=C_1x+C_2xln(x)\] So for the other part, I divided by x^2 to get that \[R(x)=\ln(x)/x^2\] Then found the Wronskian: \[W(u_1,u_2)=W(x,xln(x))=\det(\left[\begin{matrix}x & xln(x) \\ 1 & \ln(x)+1\end{matrix}\right]=xln(x)+x-xln(x)=x \neq 0 \forall x \in (0, \infty)\]

OpenStudy (anonymous):

So finally we get to the part I'm stuck on, my calculation of C_1 \[C_1=\int\limits_{}^{}\frac{ R(x)u_2 }{ W }dx= \int\limits_{}^{}\frac{ \frac{ \ln(x) }{ x^2 }xln(x) }{ x }dx=\int\limits_{}^{}\frac{ \ln(x)^2 }{ x^2}dx\]

OpenStudy (anonymous):

@zepdrix could you help with ODE's?

OpenStudy (anonymous):

oops forgot the negative above...heh heh: \[C_1=\int\limits_{}^{}-\frac{ R(x)u_2 }{ W }dx=\int\limits_{}^{}-\frac{ \ln(x)^2 }{ x^2 }dx\]

OpenStudy (anonymous):

@Directrix you've helped me before, any chance you could do so again?

OpenStudy (anonymous):

@ganeshie8 it was suggested that I tag you by mimi in the chat box.

OpenStudy (mendicant_bias):

I'd suggest integration by parts, possibly? This is more of an integration problem than anything else at the moment. Lemme take a shot.

OpenStudy (mendicant_bias):

Choosing ln(x^2) as the thing that will be derived and choosing 1/x^2 as the thing that will be integrated will eventually rid you of the 1/x^2 altogether and give you another ln(x).

OpenStudy (anonymous):

Aye, it more or less is. Integral should be set up correctly. Key being should be. Just cannot see the integration. Been working with it for about an hour now. Just cannot see it for some reason.

OpenStudy (freckles):

well it seems the way @Mendicant_Bias suggested works if I didn't make a mistake

OpenStudy (anonymous):

There is a small problem its (ln(x))^2 not ln(x^2)

OpenStudy (freckles):

and it it (ln(x))^2

OpenStudy (freckles):

it should work with (ln(x))^2 as the differentiating term and 1/x^2 as the integrating term

OpenStudy (freckles):

2 rounds of integration by parts is what I did

OpenStudy (anonymous):

Let me try that again, maybe I screwed it up somewhere the first time.

OpenStudy (anonymous):

yup just screwed up the integration. and well eliassaab, thanks but I think you did the case where the roots are imaginary not the case where they are identical. The correct answer is here http://www.wolframalpha.com/input/?i=dsolve%5Bx%5E2y%27%27-xy%27%2By%3D%3Dln%28x%29%5D

OpenStudy (anonymous):

If I could medal both, I would but I'll medal the first person to mention the correct way of doing it.

OpenStudy (anonymous):

I appreciate it. Did all the complex stuff just to screw up on the integral, lol. Thanks guys.

OpenStudy (anonymous):

I did the + x y', sorry

OpenStudy (anonymous):

No worries. I appreciate anyone who genuinely trys to help.